# Identification and Control of Neutral Anyons

### v1

@@
Within the boundary of the charge gap, drag measurements reveal an intriguing crossover. In the region near $\Delta\nu$ $=0$, the bilayer sample exhibits vanishing conductance in both parallel flow and counterflow measurements, as illustrated in Fig. 3c. As both drive and drag currents are zero, the drag ratio is ill-defined, which is indicated by the gray-shaded area in Fig. 3b and c. As $\Delta\nu$ is detuned from zero, the counterflow conductance rises from zero while the parallel flow conductance remains vanishingly small. This leads to a regime characterized by a perfect drag response, represented by dark red in the middle and right panels of panel of Fig. 3b and indicated in the red-shaded area of Fig. 3c. These transport responses point to excitonic pairing, driven by the doping-induced population of quasiparticles and quasiholes.
@@
Notably, tuning the total filling fraction, $\nu_{\text{total}}$, results in a region with a zero drag ratio, outlined by blue dashed lines and highlighted by the blue-shaded area in the middle and right panels of Fig. 3b, respectively. Fig. 3e plots the transport responses as $\nu_{1}$ is detuned from the Laughlin state. The onset of $G_{\text{PF}}$ indicates that the charge gap is suppressed by the presence of quasiparticles in layer 1 (black solid line in Fig. 3e). However, the absence of quasiholes in layer 2 prevents the formation of exciton pairs, thereby suppressing the counterflow conductance, $G_{\text{CF}}$ (blue solid line in Fig. 3e). This results in a non-zero $G_{\text{PF}}$ and a zero $G_{\text{CF}}$, confirming that exciton formation requires doping with varying $\Delta\nu$ while maintaining a constant $\nu_{\text{total}}$.
@@
Having established quasiparticle pairing in the Laughlin state, we now explore other neutral modes by examining transport properties along the equal density line, which is defined by $\nu_{1}=\nu_{2}$ and depicted by the blue horizontal line in Fig. 4a. Along this line, transport measurement uncovers three distinct types of neutral modes in the form of bilayer excitons, as shown in Fig. 4b-d, with their corresponding locations marked by green, blue, and red circles in Fig. 4a. At $\nu_{1}=\nu_{2}=1/2$ (green circle), the perfect drag response, accompanied by vanishing $G_{\text{PF}}$, points to an excitonic state described by the (111) wavefunction  Nandi2012exciton ; Eisenstein2014 ; Zeng2023solid ; Zhang2024fractionalexciton ; Li2017superfluid ; Liu2017superfluid ; Kellogg2003 ; Kellogg2004 ; Tutuc2004counterflow . These excitons consist of components with integer charge, which combine to form composite bosons (see inset in Fig. 4b, framed by a green square). As $\nu_{1}$ and $\nu_{2}$ deviate from half filling, a series of incompressible states–evidenced by vanishing $G_{\text{PF}}$–emerge at positions expected from the Jain sequence, defined by $\nu_{1}^{\ast}=\nu_{2}^{\ast}=N$, with $N\in\mathbb{Z}$.
This sequence of Jain states exhibit two interconnected hierarchical behaviors. First, the magnitude of the charge gap, reflected by the width of the insulating features in $G_{\text{PF}}$, is enhanced at lower effective filling. Second, quasiparticle pairing is suppressed at $\nu_{1}=\nu_{2}=1/3$ due to the large charge gap, but re-emerges as the charge gap decreases at higher effective filling. This is demonstrated by the perfect drag response and non-zero counterflow conductance at $\nu_{1}=\nu_{2}=2/5$, $3/7$, and $4/9$. As shown in the insets of Fig. 4b (framed by blue squares), these excitons have fractionally charged constituents that obey anyonic statistics, leading the neutral modes to exhibit anyonic behavior as well.
Most remarkably, transport measurement unveils an unexpected neutral mode at $\nu_{1}=\nu_{2}=3/8$, characterized by vanishing $G_{\text{PF}}$, enhanced counterflow conductance, $G_{\text{CF}}$, and a peak in the drag ratio approaching unity. Two observations suggest that this state is linked to the Jain sequences of FQHE where composite fermions (CFs) are constructed with two and four intralayer flux attachments, but notably without interlayer flux attachments  Li2019pairing ; Zhang2024fractionalexciton .
First, the state at $\nu_{1}=\nu_{2}=3/8$ is surrounded by these Jain sequences of FQHE states, which appears at effective integer filling of CFs. These Jain states, marked by blue, black, and gray circles in Fig. 4a, are identified by the vanishing $G_{\text{PF}}$ (dark blue) in Fig. 4f. Second, the FQHE state at $\nu_{1}=\nu_{2}=3/8$ exhibits a $B$-induced transition similar to transitions observed at $\nu_{1}=\nu_{2}=1/3$, $2/5$, and $3/7$, as marked by black horizontal lines in Fig. 4e. The low-$B$ states are described by two-component wavefunctions  Li2019pairing ; Zhang2024fractionalexciton , whereas the high-$B$ states are one-component in nature, which fall into the conventional Jain sequence.
@@
Despite its connection with the Jain sequence, the state at $\nu_{1}=\nu_{2}=3/8$ appears at a half-filled $\Lambda$ level, suggesting an exotic origin driven by residual interactions between CFs  Pan.03 ; xia2004electron ; pan2015fractional ; bellani2010optical ; samkharadze2015observation . The emergence of this state may follow a similar mechanism as the celebrated $\nu=5/2$ state. In this case, electrons in the lowest LL capture two-flux quanta and populate $\Lambda$ levels in response to the residual magnetic field. When the second $\Lambda$ level is half-filled, the two-flux composite fermions capture two additional flux quanta, undergo Cooper instability, and condensed into a $p-$wave paired state. This scenario is supported by previous numerical simulations, identifying a relative to the Bonderson-Slingerland (BS) wavefunction as the stable ground state at $\nu=3/8$ in the lowest LL of a single 2D confinement  scarola2002possible ; Mukherjee2012Pfaffian ; mukherjee2014possible ; Bonderson2008 . This describes a non-Abelian state similar to the anti-Pfaffian state at $\nu=5/2$.
@@
Intriguingly, previous studies on ultra-high mobility GaAs samples have reported the FQHE at $\nu=3/8$, originating from residual composite fermion interaction  wang2022even ; wang2023next ; wang2023fractional . However, this exotic state has been notably absent in prior experimental studies involving monolayer graphene samples with similar doubly-encapsulated geometry  Zibrov2018even ; Zeng2019 ; Chen2019 ; Polshyn2018corbino . It should be noted that LL mixing plays a significant role in facilitating interaction between composite fermions, creating favorable condition for pairing at half-filled $\Lambda$ levels  zhao2023composite ; wang2023next . Indeed, these unconventional states in GaAs were observed in hole-doped LL filling rather than electron-doped, presumably because the larger effective mass of holes, and consequently larger LL mixing, enhances interactions between CFs wang2022even ; wang2023next ; wang2023fractional . In quantum Hall graphene bilayers, we expect that the unique Coulomb environment reduces short-range repulsion in the CF sea, making it more favorable for CFs to form Cooper pairs. At large magnetic field, the absence of interlayer correlations restricts Cooper pairing to occur within each individual layer. As a result, the bilayer system offers a favorable platform to stabilize novel composite fermion states driven by intra-layer correlations, which are distinct from those observed in monolayer graphene samples.
At $\nu=3/8$, quasiparticle excitations in the sample bulk likely obeys non-Abelian quantum statistics, which arises from the non-trivial fusion rules between the distinct topological sectors. In a quantum Hall bilayer, we expect that quasiparticles at $\nu_{1}=\nu_{2}=3/8$ belong to the same topological sector. This results in a bilayer state with charge-neutral excitons that follow non-Abelian statistics, a highly unusual construction (see Methods section for more detailed discussion on the likely non-Abelian order of the bilayer state).
In summary, our findings reveal charge-neutral anyons formed through the pairing of quasiparticles and quasiholes. These novel anyons, emerging from the well-studied Jain sequence of FQHE states, highlight the charge-neutral channel as an unexplored pathway that has remained largely hidden until this work. Neutral anyons represent a new class of quantum particles distinct from charged particles. Their charge neutrality leads to a unique form of correlation beyond conventional Coulomb interaction, especially in the context of bilayer excitons. The flow dynamics of neutral anyons can be directly explored using counterflow transport, while their population can be uniquely controlled through layer-asymmetric field-effect doping. Together, these properties and tunability open a new dimension for constructing and engineering exotic quantum phases of matter, which will certainly inspire extensive future research interests.
@@ acknowledgments
This material is based on the work supported by the Air Force Office of Scientific Research under award no. FA9550-23-1-0482. N.J.Z., R.Q.N., and J.I.A.L. acknowledge support from the Air Force Office of Scientific Research. N.J.Z. acknowledges partial support from the Jun-Qi fellowship. R.Q.N. and J.I.A.L. acknowledges partial support from the National Science Foundation EPSCoR Program under NSF Award OIA-2327206. Any opinions, findings and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect those of the National Science Foundation. This work was performed in part at the Aspen Center for Physics, which is supported by a grant from the Alfred P. Sloan Foundation (G-2024-22395). A portion of this work was performed at the National High Magnetic Field Laboratory, which is supported by National Science Foundation Cooperative Agreement No. DMR-1157490 and the State of Florida. N.B. and D.E.F. were supported in part by the National Science Foundation under Grant No. DMR-2204635. K.W. and T.T. acknowledge support from the JSPS KAKENHI (Grant Numbers 21H05233 and 23H02052) and World Premier International Research Center Initiative (WPI), MEXT, Japan.
@@ II.1 Fractional charge and statistics of excitons belonging to ${}^{2}_{0}$ family
The minimal solutions $\ell_{i}$ can also be used to compute the exciton statistics. Before we compute the exciton statistics, we notice that due to the definition $\ell_{i}\equiv\sum_{m}^{|\nu_{i}^{*}|}\left[{\bf l}_{i}\right]_{m}$, there is an ambiguity in defining the excitation corresponding to the integer vector ${\bf l}_{i}$. This ambiguity suggests that there are distinct solutions of ${\bf l}$ that all correspond to the same exciton charge, or for that matter, a generic quasiparticle charge. However, to make sure all these distinct solutions describe the same particle, we need to check if all such excitations result in the same mutual and self statistical phases modulo $2\pi$. We proved that the distinct solutions indeed result in the same statistical phases in our previous report (see Ref. Zhang2024fractionalexciton ) for a general bilayer ${}^{a}_{b}$ family. This guarantees that for any choice of ${\bf l}_{i}$ corresponding to the solutions obtained $\ell_{i}$ indeed describes the same minimal exciton as with any other choice of ${\bf l}_{i}$ preserving $\ell_{i}$. We thus set $\left[{\bf l}_{1}\right]_{1}=\ell_{1}$ and $\left[{\bf l}_{2}\right]_{1}=\ell_{2}$, and all the remaining elements are set to zero $[{\bf l}_{i}]_{n}=0$ $\forall~{}n\in[2,|\nu_{i}^{*}|]$. The exciton self-statistics is computed as,

### v2

@@
Within the boundary of vanishing $G_{\text{PF}}$, drag measurements reveal the emergence of exciton pairing. Near $\Delta\nu$ $=0$, the bilayer sample exhibits vanishing conductance in both parallel flow and counterflow measurements, as illustrated in Fig. 3c. As both drive and drag currents are zero, the drag ratio is ill-defined, which is indicated by the gray-shaded area in Fig. 3b and c. As $\Delta\nu$ deviates from zero, the counterflow conductance rises from zero while the parallel flow conductance remains vanishingly small. This leads to a regime characterized by a perfect drag response, represented by dark red in the middle and right panels of panel of Fig. 3b and indicated in the red-shaded area of Fig. 3c. These transport responses point to excitonic pairing, driven by the doping-induced population of quasiparticles and quasiholes.
@@
Notably, tuning $\nu_{\text{total}}$ at $\Delta\nu$$=0$ results in a region with a zero drag ratio, outlined by blue dashed lines and highlighted by the blue-shaded area in the middle and right panels of Fig. 3b, respectively. Fig. 3e plots the transport responses as $\nu_{1}$ is detuned from the Laughlin state. The onset of $G_{\text{PF}}$ indicates that the charge gap is suppressed by the presence of quasiparticles in layer 1 (black solid line in Fig. 3e). However, the absence of quasiholes in layer 2 prevents the formation of exciton pairs, thereby suppressing the counterflow conductance, $G_{\text{counter}}$ (blue solid line in Fig. 3e). This results in a non-zero $G_{\text{PF}}$ and a zero $G_{\text{counter}}$, confirming that exciton formation requires doping with varying $\Delta\nu$ while maintaining a constant $\nu_{\text{total}}$.
@@
Having established quasiparticle pairing in the Laughlin state, we now expand our investigation to other types of excitons along the equal density line, defined by $\nu_{1}=\nu_{2}$ and depicted as the horizontal blue line in Fig. 4a. Along this line, transport measurements reveal three distinct types of excitonic states, with their respective positions marked by green, blue, and red circles in Fig. 4a. At $\nu_{1}=\nu_{2}=1/2$ (green circle), the perfect drag response accompanied by vanishing $G_{\text{PF}}$ indicates an excitonic state described by the (111) wavefunction Nandi2012exciton ; Eisenstein2014 ; Zeng2023solid ; Zhang2025fractionalexciton ; Li2017superfluid ; Liu2017superfluid ; Kellogg2003 ; Kellogg2004 ; Tutuc2004counterflow . These excitons are composed of integer-charged constituents that combine to form composite bosons (see inset in Fig. 4b, green frame).
As $\nu_{1}$ and $\nu_{2}$ deviate from half-filling, a series of FQHE states emerge along the Jain sequence, as evidenced by vanishing $G_{\text{PF}}$ at $\nu_{1}^{\ast}=\nu_{2}^{\ast}=2$, 3, 4, and 5 (Fig. 4b). At these fillings, we observe a perfect drag ratio and non-zero counterflow conductance, providing strong support for the presence of interlayer excitons. We argue that the pairing mechanism at these fillings is analogous to that at $\nu_{1}=\nu_{2}=1/3$, where excitons form via interlayer pairing of quasiparticles and quasiholes.
Notably, the interlayer excitons observed along the Jain sequence should be distinguished from those described by the (111) state at $\nu_{1}=\nu_{2}=1/2$ Eisenstein2014 ; Li2017superfluid ; Liu2017superfluid ; Liu2022crossover , as well as those associated with two-component fractional quantum Hall effect Zhang2025fractionalexciton . Ground states described by two-component wavefunctions, such as the (111) and (333) states Halperin1983 ; Wen1992 , arise in regimes where strong interlayer correlation plays a critical role in defining the FQHE phase. In contrast, the regime explored in this work is characterized by weak interlayer coupling, where FQHE states are stabilized primarily by intralayer Coulomb interactions.
The regimes of strong and weak interlayer correlation can be identified based on the hierarchical behavior of the FQHE sequence Jain.03 ; Li2019pairing . The boundary between these two regimes are revealed in Figure 4e, which plots $G_{\text{PF}}$ as a function of LL filling and $B$, measured along the equal density line. The regime of strong interlayer correlation occurs at $B<10$ T, where the FQHE sequence aligns with composite fermions incorporating interlayer flux attachment Li2019pairing ; Zhang2025fractionalexciton . In contrast, at $B>15$ T, the FQHE sequence exhibits the characteristic hierarchical structure of the Jain sequence. These two regimes are separated by a series of $B$-induced transitions at $\nu_{1}=\nu_{2}=1/3$, $2/5$, and $3/7$, marked by black horizontal lines in Fig. 4e, where the charge gap is weakened due to a transition between two-component and one-component FQHE ground states.
Amongst the Jain sequence of FQHE states in the regime of weak interlayer correlation, transport measurements unveil an unexpected FQHE state at $\nu_{1}=\nu_{2}=3/8$. This state emerges at $B=12$ T and persists beyond $B=31$ T (Fig. 4b-e), offering strong support that the underlying charge order is single-component in nature. However, $\nu_{1}=\nu_{2}=3/8$ lies outside the Jain sequence of FQHE, indicating an exotic type of order.
In the following, we propose a plausible description of relevant physics at $\nu_{1}=\nu_{2}=3/8$.
@@
We first discuss the ground state wavefunction in each graphene layer, which defines the nature of the charge gap. Since this filling corresponds to half-filled $\Lambda$ level of component fermions  Pan.03 ; xia2004electron ; pan2015fractional ; bellani2010optical ; samkharadze2015observation , the emergence of a robust FQHE state could follow a similar mechanism as the celebrated $\nu=5/2$ state. In this case, electrons in the lowest LL capture two-flux quanta and populate $\Lambda$ levels in response to the residual magnetic field. When the second $\Lambda$ level is half-filled, the two-flux composite fermions capture two additional flux quanta, undergo Cooper instability, and condensed into a $p-$wave paired state. This scenario is supported by previous numerical simulations, identifying a relative to the Bonderson-Slingerland (BS) wavefunction as the stable ground state at $\nu=3/8$ in the lowest LL of a single 2D confinement  scarola2002possible ; Mukherjee2012Pfaffian ; mukherjee2014possible ; Bonderson2008 . This describes a non-Abelian state similar to the Pfaffian state at $\nu=5/2$  Halperin2020fractional . In this scenario, quasiparticle excitations in the sample bulk likely obeys non-Abelian quantum statistics, which arises from the non-trivial fusion rules between the distinct topological sectors.
Most interestingly, the transport response at $\nu_{1}=\nu_{2}=3/8$ offers clear indication for interlayer exciton pairing. Besides vanishing $G_{\text{PF}}$, this state exhibits an enhanced counterflow conductance, as well as a peak in the drag ratio approaching unity. Given the robust charge gap and the single-component nature of the ground state, we argue that exciton pairing occurs between quasiparticle and quasihole excitations. Notably, if the FQHE state at $\nu_{1}=\nu_{2}=3/8$ is non-Abelian in nature, the resulting exciton will follow non-Abelian statistics as well, giving rise to a highly unusual construction (see Methods section for more detailed discussion on the likely non-Abelian order of the bilayer state).
@@
Previous studies on ultra-high mobility GaAs samples have reported the FQHE at $\nu=3/8$, originating from residual composite fermion interaction  wang2022even ; wang2023next ; wang2023fractional . However, this exotic state has been notably absent in prior experimental studies involving monolayer graphene samples with similar doubly-encapsulated geometry  Zibrov2018even ; Zeng2019 ; Chen2019 ; Polshyn2018corbino . It should be noted that LL mixing plays a significant role in facilitating interaction between composite fermions, creating favorable condition for pairing at half-filled $\Lambda$ levels  zhao2023composite ; wang2023next . Indeed, these unconventional states in GaAs were observed in hole-doped LL filling rather than electron-doped, presumably because the larger effective mass of holes, and consequently larger LL mixing, enhances interactions between CFs wang2022even ; wang2023next ; wang2023fractional . In quantum Hall graphene bilayers, we expect that the unique Coulomb environment reduces short-range repulsion in the CF sea, making it more favorable for CFs to form Cooper pairs. At large magnetic field, the absence of interlayer correlations restricts Cooper pairing to occur within each individual layer. As a result, the bilayer system offers a favorable platform to stabilize novel composite fermion states driven by intra-layer correlations, which are distinct from those observed in monolayer graphene samples.
In summary, we report a new type of bilayer exciton in quantum Hall graphene bilayers, constructed through pairing between quasiparticles and quasiholes of single-component FQHE states. At the Laughlin state, fractionally charged quasiparticles exhibit a braiding phase of $32\pi/$ 3. Because the particle and hole components of the exciton reside in separate layers, the resulting exciton inherits a total braiding phase of $34\pi/$ 3. This leads to a highly unusual composite particle that is overall charge-neutral but obeys anyonic quantum statistics. The flow dynamics of neutral anyons can be directly explored using counterflow transport, while their population can be uniquely controlled through layer-asymmetric field-effect doping. Together, this new type of bilayer exciton open a new frontier for engineering exotic quantum phases, poised to inspire broad future research directions.
@@ acknowledgments
This material is based on the work supported by the Air Force Office of Scientific Research under award no. FA9550-23-1-0482. N.J.Z., R.Q.N., and J.I.A.L. acknowledge support from the Air Force Office of Scientific Research. N.J.Z. acknowledges partial support from the Jun-Qi fellowship. R.Q.N. and J.I.A.L. acknowledges partial support from the National Science Foundation EPSCoR Program under NSF Award OIA-2327206. Any opinions, findings and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect those of the National Science Foundation. This work was performed in part at the Aspen Center for Physics, which is supported by a grant from the Alfred P. Sloan Foundation (G-2024-22395). A portion of this work was performed at the National High Magnetic Field Laboratory, which is supported by National Science Foundation Cooperative Agreement No. DMR-1157490 and the State of Florida. N.B. and D.E.F. were supported in part by the National Science Foundation under Grant No. DMR-2204635. K.W. and T.T. acknowledge support from the JSPS KAKENHI (Grant Numbers 21H05233 and 23H02052) and World Premier International Research Center Initiative (WPI), MEXT, Japan. Part of this work was enabled by the use of pyscan (github.com/sandialabs/pyscan), scientific measurement software made available by the Center for Integrated Nanotechnologies, an Office of Science User Facility operated for the U.S. Department of Energy.
@@ II.1 Fractional charge and statistics of excitons belonging to ${}^{2}_{0}$ family
The minimal solutions $\ell_{i}$ can also be used to compute the exciton statistics. Before we compute the exciton statistics, we notice that due to the definition $\ell_{i}\equiv\sum_{m}^{|\nu_{i}^{*}|}\left[{\bf l}_{i}\right]_{m}$, there is an ambiguity in defining the excitation corresponding to the integer vector ${\bf l}_{i}$. This ambiguity suggests that there are distinct solutions of ${\bf l}$ that all correspond to the same exciton charge, or for that matter, a generic quasiparticle charge. However, to make sure all these distinct solutions describe the same particle, we need to check if all such excitations result in the same mutual and self statistical phases modulo $2\pi$. We proved that the distinct solutions indeed result in the same statistical phases in our previous report (see Ref. Zhang2025fractionalexciton ) for a general bilayer ${}^{a}_{b}$ family. This guarantees that for any choice of ${\bf l}_{i}$ corresponding to the solutions obtained $\ell_{i}$ indeed describes the same minimal exciton as with any other choice of ${\bf l}_{i}$ preserving $\ell_{i}$. We thus set $\left[{\bf l}_{1}\right]_{1}=\ell_{1}$ and $\left[{\bf l}_{2}\right]_{1}=\ell_{2}$, and all the remaining elements are set to zero $[{\bf l}_{i}]_{n}=0$ $\forall~{}n\in[2,|\nu_{i}^{*}|]$. The exciton self-statistics is computed as,
@@
Figure S1: Activated behavior of quasiparticle excitations. (a) Arrhenius plot of bulk conductance measured in the parallel flow $G_{\text{PF}}$ (black) and counterflow $G_{\text{counter}}$ (blue) configurations at $\nu_{1}=\nu_{2}=1/3$ and $B$ = 28 T. The activated behavior of $G_{\text{PF}}$ is directly linked to the charge gap, $\Delta_{\text{charge}}$, which represents the energy required to create thermally excited quasiparticles and quasiholes. Interestingly, the activated behavior of $G_{\text{counter}}$ reveals the energy required to generate a charge-neutral excitation, $\Delta_{\text{exciton}}$, which closely aligns with $\Delta_{\text{charge}}$. This provides direct evidence that both quasiparticles and the neutral mode are generated through the same thermal process.